Warriorz13:

What is the half-life of 99Tc if a 500 g sample decays to 62.5 g in 639,000 years?

Warriorz13:

@JustSaiyan

JustSaiyan:

500 g ÷ 8 = 62.5 g 1/2^3 = 1/8 639,000 is 3 half-lifes 639,000 ÷ 3 =213000 half-life is 213000 years

JustSaiyan:

Got it?

Warriorz13:

Yea.. I'll probably still post the. Tho, also doing nuclear reactions.